OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10? A. 605 B. 660 C. 748 D. 816 Please don't just give me the answer, I want to know how to do it. I can't figure it out.

OpenStudy (cwrw238):

12th term = a1 + 11d where a1 = first term and d = common difference -10 so we have 13 = a1 -11*-10 solve for a1 then find the sum using the formula Sn = (n/2) [ a1 + L] where L = last term

OpenStudy (anonymous):

Use the explicit formula to find \[a _{1}\] then use the formula below to find the sum \[S _{n} = n(\frac{ a _{1}+a _{n} }{ 2 })\]

OpenStudy (cwrw238):

nth term of an AS = a1 + (n - 1)d

OpenStudy (anonymous):

So what is L?

OpenStudy (cwrw238):

L = 13 as given

OpenStudy (anonymous):

I overlooked that. I apologize.

OpenStudy (cwrw238):

thats ok

OpenStudy (anonymous):

So we have 13 = a1 -11*-10. Do i subtract 11?

OpenStudy (anonymous):

or add*

OpenStudy (cwrw238):

no do the multiplication first

OpenStudy (cwrw238):

13 = a1 + 110

OpenStudy (anonymous):

oh okay. so 13=a1+110? since it was two negatives?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

okay then subtract 110 from 13 right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

a1=-97

OpenStudy (cwrw238):

right now use the formula for sum of n terms

OpenStudy (cwrw238):

n = number of terms in the As which is 12

OpenStudy (anonymous):

Sn = (n/2) [ a1 + L] so S(12)= (12/2) (-97+13) ?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

S12=(6)(-84)

OpenStudy (cwrw238):

sujuex's formula is same as mine but written in different way

OpenStudy (anonymous):

It looked about the same but just more confusing kind of.

OpenStudy (anonymous):

I got -504 but that isn't one of the choices. :(

OpenStudy (cwrw238):

you wont get one of the options as an answer I know Maybe you havent go the question right

OpenStudy (cwrw238):

maybe d = 10 ?

OpenStudy (anonymous):

it's from Fvls so i don't know. :/

OpenStudy (anonymous):

i could try that.

OpenStudy (anonymous):

the equation that is in my lesson is Sn =  [2a1 + (n – 1)d] could i try to plug them in here?

OpenStudy (cwrw238):

sorry - thers a mistake

OpenStudy (cwrw238):

13 = a1 +1*-10 - not a double minus

OpenStudy (anonymous):

ohhh okay. 13=a1+(-10)

OpenStudy (cwrw238):

so 13 = a1 - 11*10 and a1 = 123

OpenStudy (anonymous):

okay so S12 = (12/2) [ 123 + 13]

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

6*136=816

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

Yay! thank you so much. And i am so sorry to ask this but could you guide me through another? i think i've got it down but could you tell me if i got it right?

OpenStudy (cwrw238):

ok I've got 10 minutes then i must go

OpenStudy (anonymous):

What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374? A. 11 B. 11.5 C. 12 D. 12.5

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

would i use the same equation?

OpenStudy (cwrw238):

you know the sum of n terms and n = 43 and you also know a1 so use the formula Sn = (n/2)[2a1 + (n - 1)d] 9374 = (43/2) [2*-13 + (43-1)d] 9374 = 21.5[ -26 + 42d] solve this to find value of d

OpenStudy (anonymous):

would i add -26 and 42?

OpenStudy (anonymous):

or subtract 21.5 from 9374? if so 9352.5=(-26+42d) right?

OpenStudy (cwrw238):

no - remember when you have a mixture of adds and multiplies you do the multiply first. In any case you cannot add unlike terms . 26 and 42d are unlike terms. in this case you have the brackets so you multiply first the -26 then the + 42d by 21.5 use ur calculator. then solving for d is easy sorry gotta go

OpenStudy (anonymous):

oh okay. uhm 9374= -559 + 903d

OpenStudy (anonymous):

9374= -559 + 903d add 559 divide by 903 and the answer is 11. THANK YOU! I could just hug you. I know you're gone but yo're wonderful thanks so much.